OpenStudy (anonymous):

plz help i will medal Solve the word problem. The base of a triangle with height h and area A is given by the formula b = 2A/h What is the base of a triangle with area 17 cm2 and height 5 cm? A. 3.4 cm B. 3.8 cm C. 6.8 cm D. 11 cm

OpenStudy (ja1):

First plug in the numerical values

OpenStudy (luigi0210):

Omg, you're poop xD

OpenStudy (ja1):

xD sowwy luigi :3

OpenStudy (anonymous):

wat would that be ?

OpenStudy (ja1):

\[b = 2\frac{ 17 }{ 5 }\]

OpenStudy (anonymous):

ook

OpenStudy (ja1):

So can you solve that?

OpenStudy (anonymous):

do i need to div?

OpenStudy (ja1):

yes a fraction means divide :)

OpenStudy (anonymous):

\[2\frac{ 17 }{ 5 }\]

OpenStudy (anonymous):

is ....

OpenStudy (anonymous):

wait how would i div \[2\frac{ 17 }{ 5 }\]

OpenStudy (anonymous):

that make no scene to me

OpenStudy (ja1):

Think of it like this:\[2(\frac{ 17 }{ 5 })\]

OpenStudy (ja1):

You need to divide the 17/5 first

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i go 3.4

OpenStudy (ja1):

Did you get the answer @Jamielynne ?

OpenStudy (ja1):

Perfect, so now you do:\[2(3.4)\]

OpenStudy (anonymous):

yes i did @JA1

OpenStudy (anonymous):

uummmmmm.....

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

uum let me c/think

OpenStudy (ja1):

sorry i crashed, anyways did you get it?

OpenStudy (anonymous):

wat i dont get wat u said?

OpenStudy (ja1):

My internet crashed,

OpenStudy (anonymous):

o mine to

OpenStudy (ja1):

ya the serve must be at peak, anyways did you get what 2(3.4) is?

OpenStudy (anonymous):

yes ty

OpenStudy (ja1):

oh ok np glad to help :)

OpenStudy (anonymous):

I have to disagree with you because to find the area of a triangle you have to find the base and height and divide by 2 or multiply by 1/2 and if the base is 3.4 like you previously stated it would be 3.4*5=17 and then you would have to divide by 2 or multiply by 1/2 and you would get 8.5cm2 as the area. As you can see this wouldn't match the area given to you. This is why I disagree but since this is a year old it wouldn't matter if I explained to you what the answer is and it would help other viewers to understand as well. So the correct answer is 6.8. I say this because you focused on it being half and half of 6.8 is 3.4 but that 3.4 is not correct. Someone can simply put 6.8 because I said so and it makes since but you should understand how I got this answer. So 6.8*5=34 and 34/2 or 34*1/2 is 17 which would give you the triangle's area of 17cm2. So as proven the answer is actually C)6.8. Brought to you by a fangirl of Heroes of Olympus.